OpenStudy (anonymous):

perform the indicated operation: 5/2x+3y ÷ 10/ 4x^2-9y^2

OpenStudy (anonymous):

Is it 5 10 ------- divided by ---------- 2x+3y 4x^2-9y^2 or 5 10 ---- + 3y divided by ------ - 9y^2 2x 4x^2

OpenStudy (anonymous):

the first one.

OpenStudy (anonymous):

\[=(\frac{5}{2x+3y})(\frac{4x^2-9y^2}{10})\]This becomes:\[(\frac{1}{2x+3y})(\frac{(2x+3y)(2x-3y)}{5})\]Cancel common factor of 2x+3y:\[\frac{2x-3y}{5}\]

OpenStudy (anonymous):

5 10 ------- divided by ---------- 2x+3y 4x^2-9y^2 5 4x^2-9y^2 ------- * ----------- 2x+3y 10 5 [(2x)² - (3y)²] 5 [(2x + 3y) (2x - 3y)] 2x - 3y --------------- = --------------------- = ----------- 10 (2x + 3y) 10 (2x + 3y) 2

OpenStudy (anonymous):

eseidi has made a slight mistake.....

OpenStudy (anonymous):

oooops...yeah the denominator is 2 not 5 as I said lol sorry